In a group of 102 employees, there are three times as many employees on the day shift as on the
night shift, and two more on the swing shift than on the night shift. How many are on each shift?

Answer by ikleyn(52781)

.

The problem says

    x + 3x + (x+2) = 102,  where x is the number of the night employees.


From it,

    5x = 102 - 2 = 100,

     x           = 100/5 = 20.


ANSWER.  Night shift: 20 employees;  day shift: 3*20 = 60 employees;  swing shift: 20+2 = 22 employees.

Solved.